OpenStudy (anonymous):

Question: Part A: Divide (9x^4y^3 + 3x^3y^2 - 6x^2y - 12x^2y^4) by -3x^2y. Show your work, and justify each step. Part B: How would your answer in Part A be affected if the x2 variable in the denominator was just an x? Part C: What is the degree and classification of the polynomial you got in Part A?

OpenStudy (tkhunny):

You could examine one numerator term at a time. Let's see the first one. \(\dfrac{9x^4y^3}{-3x^2y}\)

OpenStudy (anonymous):

So I have to divide 9x^4y^3 by -3x^2y ?

OpenStudy (anonymous):

The first think I think you need to examine is this simple example \[\frac{ 10 }{ 5*2 } \] can you break this down somehow?

OpenStudy (tkhunny):

Perhaps the word "divide" is scary? Don't say it. Just simplify the fraction as it has been presented. \(\dfrac{9x^{4}y^{3}}{−3x^{2}y} = \dfrac{9}{-3}\cdot\dfrac{x^{4}}{x^{2}}\cdot\dfrac{y^{3}}{y}\) Does that look any better?

OpenStudy (anonymous):

@tkhunny Sooo it'd be -3x^6y^4 :)

OpenStudy (anonymous):

@knivez I'm confused... :/

OpenStudy (anonymous):

you can do \[ \frac{ 10 }{ 5 } = 2\] and then do \[\frac{2}{2} = 1 \] and then you can apply this same method to your problem.

OpenStudy (anonymous):

Ohh okay :) Thank you.

OpenStudy (anonymous):

so first divide by \[-3x^{2}\] and then divide by y

OpenStudy (tkhunny):

Why would you ADD the exponents in a division problem? \(x^{2}\cdot x^{4} = x^{2+4} = x^{6}\) \(\dfrac{x^{4}}{x^{2}} = x^{4 - 2} = x^{2}\)

OpenStudy (anonymous):